OpenStudy (anonymous):

Mr. Doodle’s grade distribution over the past 3 years for a course in college algebra is shown in the chart below. Grade Number A 45 B 180 C 210 D 95 F 65 I 5 If Jane plans to take a college algebra course with Mr. Doodle, determine the empirical probability that she receives an F. Mr. Doodle’s grade distribution over the past 3 years for a course in college algebra is shown in the chart below.Grade Number A 45 B 180 C 210 D 95 F 65 I 5 A=45 b=180 c=210 d=95 f=65 i=5

OpenStudy (anonymous):

13/120

OpenStudy (anonymous):

this is the correct question, sorry for the repeat

OpenStudy (amistre64):

empirical means we take what is given instead of trying to interpolate the data

OpenStudy (amistre64):

#of Fs -------------- total # of grades

OpenStudy (amistre64):

in this case, if we go off your document 65/sum of it all

OpenStudy (anonymous):

i thought u add all the number and divide by the number of the letter you looking for which is F

OpenStudy (amistre64):

almost; you got the order flipped about. probability is defined as:\[\frac{\#.desired.outcomes}{\#.total.outcomes}\]

OpenStudy (amistre64):

\[\frac{65}{sum.of.all}=\text{ empirical prob.}\]

OpenStudy (anonymous):

it didnt show explanation

OpenStudy (amistre64):

it dint show MY explanation? or it dint show an explanation in the material you have?

OpenStudy (anonymous):

65/600=.1083333333

OpenStudy (amistre64):

# of Fs -------------- = emp. prob. total # of grades

OpenStudy (amistre64):

if that 600 is good, then thats what it would amount to yes

OpenStudy (anonymous):

it total is 600

OpenStudy (amistre64):

then the prob tends to truncate to 4 decis .1083 looks good to me